# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v3

@@ 18 Ian McEwan (✓) 36 Percival (✓)Everett
style In_Context GPT-4o (In-Context) Expert 0.320 [0.214, 0.449]
style In_Context GPT-4o (In-Context) Lay 0.472 [0.407, 0.538]
style In_Context Claude (In-Context) Expert 0.380 [0.303, 0.464]
style In_Context Claude (In-Context) Lay 0.528 [0.464, 0.591]
style In_Context Gemini (In-Context) Expert 0.167 [0.097, 0.271]
style In_Context Gemini (In-Context) Lay 0.444 [0.374, 0.516]
style Fine_tuned GPT-4o (Fine-tuned) Expert 0.741 [0.684, 0.790]
style Fine_tuned GPT-4o (Fine-tuned) Lay 0.742 [0.694, 0.785]
quality In_Context GPT-4o (In-Context) Expert 0.193 [0.112, 0.313]
quality In_Context GPT-4o (In-Context) Lay 0.608 [0.541, 0.671]
quality In_Context Claude (In-Context) Expert 0.433 [0.329, 0.544]
quality In_Context Claude (In-Context) Lay 0.564 [0.499, 0.627]
quality In_Context Gemini (In-Context) Expert 0.193 [0.124, 0.289]
quality In_Context Gemini (In-Context) Lay 0.492 [0.422, 0.563]
quality Fine_tuned GPT-4o (Fine-tuned) Expert 0.578 [0.519, 0.635]
quality Fine_tuned GPT-4o (Fine-tuned) Lay 0.609 [0.562, 0.654]
@@
Tables 11 and 12 examine the writer type × reader type interactions in detail. The significant interactions in in-context prompting models (all $p<0.001$) quantify lay readers’ greater tolerance for AI-generated text. The absence of significant interactions in fine-tuned models ($p>0.4$) indicates that fine-tuning produces text that both expert and lay readers find equally compelling.
Table 11: Individual interaction terms showing differential preferences between expert and lay readers. Large positive coefficients in in-context prompting models indicate lay readers are more favorable to AI than experts.
@@
style In_Context GPT-4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
style In_Context Claude_baseline:reader_typeLay 1.428 0.336 4.249 2.15e-05
style In_Context Gemini_baseline:reader_typeLay 2.211 0.475 4.650 3.32e-06
style Fine_tuned GPT-4o_finetuned:reader_typeLay 0.015 0.372 0.042 0.967
quality In_Context GPT-4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
quality In_Context Claude_baseline:reader_typeLay 1.723 0.420 4.105 4.04e-05
quality In_Context Gemini_baseline:reader_typeLay 2.594 0.473 5.480 4.25e-08
quality Fine_tuned GPT-4o_finetuned:reader_typeLay 0.258 0.314 0.821 0.412
Table 12: Joint Wald tests for writer type × reader type interactions. Significant interactions in in-context prompting models become non-significant after fine-tuning, indicating convergence of expert and lay preferences.
@@
style In_Context writer_type:reader_type 3 24.923 1.60e-05
style Fine_tuned writer_type:reader_type 1 0.0017 0.967
quality In_Context writer_type:reader_type 3 37.584 3.46e-08
quality Fine_tuned writer_type:reader_type 1 0.674 0.412
The results in this section provide compelling statistical evidence for the paper’s central finding: fine-tuning on author-specific corpora fundamentally transforms AI-generated text from clearly inferior (as judged by experts) to preferred over human writing. The convergence of expert and lay preferences after fine-tuning suggests that the improvements are not merely superficial but represent genuine advances in literary quality and stylistic fidelity.
@@
Table LABEL:tab:s7.1 presents author-level AI preference rates using Jeffreys prior estimates to stabilize small-sample authors. The results reveal striking heterogeneity. For style, AI preference rates range from 18% (Tony Tulathimutte) to 98% (Roxane Gay), with 27 of 30 authors showing AI superiority (rate >> 0.5). For quality, the range spans 30% (Ian McEwan) to 86% (Cheryl Strayed), with 23 of 30 authors favoring AI.
@@
Table 13: Per-author AI preference rates (Jeffreys-smoothed and raw) ranked by performance. Values above 0.5 indicate AI preference. Note the wide variation across authors and the divergence between style and quality rankings for some authors.
@@
quality Cheryl Strayed 1.0 0.86 0.14 0.875
quality Marilynne Robinson 2.0 0.78 0.22 0.792
quality Colson Whitehead 3.0 0.74 0.26 0.750
quality Han Kang 4.0 0.74 0.26 0.750
quality Haruki Murakami 5.0 0.74 0.26 0.750
quality Junot Diaz 6.0 0.74 0.26 0.750
quality Rachel Cusk 7.0 0.74 0.26 0.750
quality Salman Rushdie 8.0 0.74 0.26 0.750
quality Sigrid Nunez 9.0 0.74 0.26 0.750
quality Orhan Pamuk 10.0 0.70 0.30 0.708
quality Lydia Davis 11.0 0.66 0.34 0.667
quality Percival Everett 12.0 0.66 0.34 0.667
quality Jonathan Franzen 13.0 0.62 0.38 0.625
quality Louise Erdrich 14.0 0.62 0.38 0.625
quality Annie Proulx 15.0 0.58 0.42 0.583
quality George Saunders 16.0 0.58 0.42 0.583
quality Zadie Smith 17.0 0.58 0.42 0.583
quality Chimamanda Ngozi Adichie 18.0 0.54 0.46 0.542
quality Jhumpa Lahiri 19.0 0.54 0.46 0.542
quality Margaret Atwood 20.0 0.54 0.46 0.542
quality Roxane Gay 21.0 0.54 0.46 0.542
quality Sally Rooney 22.0 0.54 0.46 0.542
quality Tony Tulathimutte 23.0 0.54 0.46 0.542
quality Min Jin Lee 24.0 0.46 0.54 0.458
quality Annie Ernaux 25.0 0.42 0.58 0.417
quality Ben Lerner 26.0 0.42 0.58 0.417
quality Kazuo Ishiguro 27.0 0.42 0.58 0.417
quality Ottessa Moshfegh 28.0 0.38 0.62 0.375
quality Yoko Ogawa 29.0 0.34 0.66 0.333
quality Ian McEwan 30.0 0.30 0.70 0.292
style Roxane Gay 1.0 0.98 0.02 1.000
style Chimamanda Ngozi Adichie 2.0 0.94 0.06 0.958
style Han Kang 3.0 0.94 0.06 0.958
style Margaret Atwood 4.0 0.94 0.06 0.958
style Ben Lerner 5.0 0.90 0.10 0.917
style Junot Diaz 6.0 0.90 0.10 0.917
style Marilynne Robinson 7.0 0.90 0.10 0.917
style Kazuo Ishiguro 8.0 0.86 0.14 0.875
style Lydia Davis 9.0 0.86 0.14 0.875
style Orhan Pamuk 10.0 0.82 0.18 0.833
style Cheryl Strayed 11.0 0.78 0.22 0.792
style Min Jin Lee 12.0 0.78 0.22 0.792
style Sigrid Nunez 13.0 0.78 0.22 0.792
style Colson Whitehead 14.0 0.74 0.26 0.750
style Haruki Murakami 15.0 0.74 0.26 0.750
style Ian McEwan 16.0 0.74 0.26 0.750
style Rachel Cusk 17.0 0.74 0.26 0.750
style Sally Rooney 18.0 0.74 0.26 0.750
style Percival Everett 19.0 0.70 0.30 0.708
style Jhumpa Lahiri 20.0 0.66 0.34 0.667
style Ottessa Moshfegh 21.0 0.66 0.34 0.667
style Zadie Smith 22.0 0.66 0.34 0.667
style Annie Proulx 23.0 0.62 0.38 0.625
style George Saunders 24.0 0.62 0.38 0.625
style Jonathan Franzen 25.0 0.62 0.38 0.625
style Salman Rushdie 26.0 0.62 0.38 0.625
style Yoko Ogawa 27.0 0.62 0.38 0.625
style Louise Erdrich 28.0 0.50 0.50 0.500
style Annie Ernaux 29.0 0.42 0.58 0.417
style Tony Tulathimutte 30.0 0.18 0.82 0.167
@@
quality 0.0070 -0.0174 0.0314 0.570 0.458 0.681 0.0122 30
style 0.0032 -0.0262 0.0326 0.729 0.595 0.863 0.0017 30
Non-parametric check. Spearman $\rho=0.138$ ($p=0.467$) for quality and $\rho=-0.140$ ($p=0.462$) for style.
@@
Table LABEL:tab:s8.1 presents the full model examining how AI detectability (Pangram score) influences preferences. The key finding is the significant interaction between pangram_score and setting (quality: $\hat{\beta}\approx 2.90$, $p<0.001$), indicating that fine-tuning reverses and substantially attenuates the negative association between detectability and preference observed in in-context prompting.
Table 15: Pangram GLM coefficients with CR2 robust standard errors (clustered by reader). Negative coefficients for pangram_score in in-context prompting indicate detection reduces preference; positive interactions with setting show that fine-tuning reverses and attenuates this penalty.
@@
quality 3840 159 (Intercept) 0.990 0.169 5.87 0.000 2.69 1.93 3.74
quality 3840 159 pangram_score -2.010 0.333 -6.03 0.000 0.13 0.07 0.26
quality 3840 159 setting (Fine-tuned) -1.020 0.175 -5.86 0.000 0.36 0.26 0.51
quality 3840 159 reader_type (Lay) -1.220 0.190 -6.42 0.000 0.30 0.20 0.43
quality 3840 159 pangram_score $\times$ setting 2.900 0.876 3.32 0.001 18.20 3.28 102.00
quality 3840 159 pangram_score $\times$ reader_type 2.480 0.377 6.56 0.000 11.90 5.67 24.90
quality 3840 159 setting $\times$ reader_type 1.250 0.198 6.33 0.000 3.50 2.37 5.15
quality 3840 159 pangram_score $\times$ setting $\times$ reader_type -3.290 1.030 -3.21 0.001 0.04 0.00 0.28
style 3840 159 (Intercept) 0.909 0.146 6.23 0.000 2.48 1.86 3.31
style 3840 159 pangram_score -1.850 0.292 -6.32 0.000 0.16 0.09 0.28
style 3840 159 setting (Fine-tuned) -0.938 0.158 -5.95 0.000 0.39 0.29 0.53
style 3840 159 reader_type (Lay) -0.832 0.167 -4.97 0.000 0.44 0.31 0.60
style 3840 159 pangram_score $\times$ setting 2.560 0.810 3.15 0.002 12.90 2.63 63.00
style 3840 159 pangram_score $\times$ reader_type 1.690 0.336 5.03 0.000 5.41 2.80 10.50
style 3840 159 setting $\times$ reader_type 0.840 0.178 4.73 0.000 2.32 1.64 3.28
style 3840 159 pangram_score $\times$ setting $\times$ reader_type -1.900 0.922 -2.06 0.039 0.15 0.02 0.91
Notes. CR2 robust standard errors clustered by reader. Readers: 28 Expert and 131 Lay (total $N_{\text{readers}}=159$).
@@
Tables 16 and 17 show model-predicted probabilities at different AI detection levels, corresponding to Figure 2F. For expert readers evaluating quality, high detection scores (0.9) reduce AI preference to $\sim$31% in in-context prompting but have a point estimate of $\sim$68% after fine-tuning (95% CI 0.370.370.890.89), demonstrating that fine-tuning makes outputs robust to detection-based skepticism.
Table 16: Predicted AI preference probabilities for style at different Pangram detection scores. Fine-tuning reverses and attenuates the relationship with detection score (with wider uncertainty at high detection).
@@
0.1 In-context Prompting Expert 0.674 0.621 0.722
0.5 In-context Prompting Expert 0.497 0.494 0.500
0.9 In-context Prompting Expert 0.320 0.273 0.372
0.1 Fine-tuned Expert 0.511 0.491 0.530
0.5 Fine-tuned Expert 0.581 0.435 0.714
0.9 Fine-tuned Expert 0.648 0.380 0.847
0.1 In-context Prompting Lay 0.515 0.483 0.547
0.5 In-context Prompting Lay 0.500 0.499 0.500
0.9 In-context Prompting Lay 0.484 0.451 0.517
0.1 Fine-tuned Lay 0.507 0.495 0.520
0.5 Fine-tuned Lay 0.557 0.461 0.649
0.9 Fine-tuned Lay 0.606 0.428 0.759
Table 17: Predicted AI preference probabilities for quality. Expert readers show strong detection sensitivity in in-context prompting ($\sim$0.69 $\to$ $\sim$0.31) that reverses after fine-tuning (point estimate at 0.9 $\sim$0.68; 95% CI 0.37–0.89).
@@
0.1 In-context Prompting Expert 0.688 0.628 0.742
0.5 In-context Prompting Expert 0.496 0.493 0.499
0.9 In-context Prompting Expert 0.306 0.254 0.363
0.1 Fine-tuned Expert 0.514 0.490 0.537
0.5 Fine-tuned Expert 0.602 0.431 0.751
0.9 Fine-tuned Expert 0.683 0.373 0.887
0.1 In-context Prompting Lay 0.454 0.421 0.488
0.5 In-context Prompting Lay 0.501 0.500 0.501
0.9 In-context Prompting Lay 0.547 0.512 0.582
0.1 Fine-tuned Lay 0.501 0.487 0.515
0.5 Fine-tuned Lay 0.509 0.403 0.614
0.9 Fine-tuned Lay 0.517 0.324 0.705
@@
Table 18 shows that cliché density has the strongest correlation with AI detection (Pearson $r=0.60$), while readability ease is negatively correlated (r=0.23r=-0.23), suggesting AI text is marked by formulaic phrases but simpler syntax. In in-context prompting conditions, approximately 16% of the detection effect on preference is mediated through cliché density; after fine-tuning, this mediation drops to a statistically insignificant 1.3%, indicating that fine-tuning substantially reduces rather than merely masking these stylistic signatures.
Table 18: Correlations between Pangram AI detection scores and stylometric features.
@@
Table 19: Cost comparison for generating 100,000 words. Fine-tuning plus inference costs (\$25–\$276) represent less than 1% of professional writer compensation (\$25,000).
@@
Note. Inference cost of \$3 per 100k words follows fig4c_cost_summary.csv; per-token assumptions are documented in the repository.
@@
    • Expert readers: 28 (planned:  25)
    • Lay readers: 131 (planned: 120)

### v4

@@ 18 Ian McEwan (✓) 36 Percival Everett (✓)
style In_Context GPT-4o (In-Context) MFA-trained 0.320 [0.21, 0.45]
style In_Context GPT-4o (In-Context) College-educated 0.525 [0.49, 0.56]
style In_Context Claude (In-Context) MFA-trained 0.380 [0.30, 0.46]
style In_Context Claude (In-Context) College-educated 0.541 [0.51, 0.57]
style In_Context Gemini (In-Context) MFA-trained 0.167 [0.10, 0.27]
style In_Context Gemini (In-Context) College-educated 0.455 [0.42, 0.49]
style Fine_tuned GPT-4o (Fine-tuned) MFA-trained 0.741 [0.68, 0.79]
style Fine_tuned GPT-4o (Fine-tuned) College-educated 0.803 [0.78, 0.82]
quality In_Context GPT-4o (In-Context) MFA-trained 0.193 [0.11, 0.31]
quality In_Context GPT-4o (In-Context) College-educated 0.621 [0.59, 0.65]
quality In_Context Claude (In-Context) MFA-trained 0.433 [0.33, 0.54]
quality In_Context Claude (In-Context) College-educated 0.597 [0.56, 0.63]
quality In_Context Gemini (In-Context) MFA-trained 0.193 [0.12, 0.29]
quality In_Context Gemini (In-Context) College-educated 0.503 [0.47, 0.54]
quality Fine_tuned GPT-4o (Fine-tuned) MFA-trained 0.578 [0.52, 0.64]
quality Fine_tuned GPT-4o (Fine-tuned) College-educated 0.699 [0.68, 0.72]
@@
Tables 11 and 12 examine the writer-type $\times$ reader-type interactions in detail. The significant interactions in in-context prompting models (all $p<0.001$) quantify college-educated general readers’ greater tolerance for AI-generated text. Unlike the previous analysis with a smaller college-educated sample, the interactions in fine-tuned models are now also significant (style $p=0.021$; quality $p=6.41\times 10^{-5}$), indicating that college-educated general readers favor fine-tuned AI by a wider margin than MFA-trained readers rather than converging to similar preference levels.
Table 11: Individual interaction terms showing differential preferences between MFA-trained and college-educated general readers. Large positive coefficients indicate college-educated general readers are more favorable to AI. In fine-tuned models, interactions are significant, indicating that reader-group differences persist after fine-tuning.
@@
style In_Context GPT-4o_baseline:reader_typeCollege-educated 1.784 0.414 4.307 1.66e-05
style In_Context Claude_baseline:reader_typeCollege-educated 1.583 0.292 5.412 6.22e-08
style In_Context Gemini_baseline:reader_typeCollege-educated 2.357 0.438 5.384 7.27e-08
style Fine_tuned GPT-4o_finetuned:reader_typeCollege-educated 0.713 0.309 2.308 0.021
quality In_Context GPT-4o_baseline:reader_typeCollege-educated 3.197 0.484 6.601 4.08e-11
quality In_Context Claude_baseline:reader_typeCollege-educated 1.935 0.381 5.073 3.91e-07
quality In_Context Gemini_baseline:reader_typeCollege-educated 2.716 0.433 6.268 3.65e-10
quality Fine_tuned GPT-4o_finetuned:reader_typeCollege-educated 1.062 0.266 3.997 6.41e-05
Table 12: Joint Wald tests for writer-type $\times$ reader-type interactions. Interactions are significant in both in-context prompting and fine-tuned models, indicating that college-educated general readers are more favorable to AI than MFA-trained readers in both conditions.
@@
style In_Context writer_type:reader_type 3 36.606 5.57e-08
style Fine_tuned writer_type:reader_type 1 5.328 0.021
quality In_Context writer_type:reader_type 3 46.884 3.68e-10
quality Fine_tuned writer_type:reader_type 1 15.978 6.41e-05
The results in this section provide compelling statistical evidence for the paper’s central finding: fine-tuning on author-specific corpora fundamentally transforms AI-generated text from clearly inferior (as judged by experts) to preferred over human writing. Rather than converging after fine-tuning, both reader groups prefer fine-tuned AI, but college-educated general readers do so by a significantly wider margin than MFA-trained readers. This pattern suggests that fine-tuning produces improvements that are meaningful to both audiences while leaving a persistent difference in how strongly each audience favors the AI output.
@@
Table LABEL:tab:s7.1 presents author-level AI preference rates using Jeffreys prior estimates to stabilize small-sample authors. Each author now has 72 pooled trials (9 from MFA-trained readers and 63 from college-educated general readers), so the pooled rates are weighted approximately 87.5% toward the college-educated group. The results reveal striking heterogeneity. For style, AI preference rates range from 18% (Tony Tulathimutte) to 98% (Roxane Gay), with 29 of 30 authors showing AI superiority (rate >> 0.5). For quality, the range spans 49% (Kazuo Ishiguro and Min Jin Lee) to 92% (Cheryl Strayed), with 28 of 30 authors favoring AI.
@@
Table 13: Per-author AI preference rates (Jeffreys-smoothed and raw) ranked by performance. Values above 0.5 indicate AI preference. Each author has 72 pooled trials (9 MFA-trained, 63 college-educated general readers). Note the wide variation across authors and the divergence between style and quality rankings for some authors.
@@
quality Cheryl Strayed 1 0.92 0.08 0.931
quality Marilynne Robinson 2 0.91 0.09 0.917
quality Han Kang 3 0.88 0.12 0.889
quality Salman Rushdie 4 0.84 0.16 0.847
quality Lydia Davis 5 0.80 0.20 0.806
quality Zadie Smith 6 0.80 0.20 0.806
quality Haruki Murakami 7 0.77 0.23 0.778
quality Junot Diaz 8 0.77 0.23 0.778
quality Orhan Pamuk 9 0.77 0.23 0.778
quality Sigrid Nunez 10 0.76 0.24 0.764
quality Colson Whitehead 11 0.73 0.27 0.736
quality Roxane Gay 12 0.71 0.29 0.708
quality Percival Everett 13 0.69 0.31 0.694
quality Jhumpa Lahiri 14 0.68 0.32 0.681
quality Rachel Cusk 15 0.68 0.32 0.681
quality Ben Lerner 16 0.66 0.34 0.667
quality Margaret Atwood 17 0.66 0.34 0.667
quality Louise Erdrich 18 0.65 0.35 0.653
quality George Saunders 19 0.64 0.36 0.639
quality Jonathan Franzen 20 0.64 0.36 0.639
quality Sally Rooney 21 0.64 0.36 0.639
quality Annie Proulx 22 0.60 0.40 0.597
quality Chimamanda Ngozi Adichie 23 0.58 0.42 0.583
quality Annie Ernaux 24 0.55 0.45 0.556
quality Tony Tulathimutte 25 0.55 0.45 0.556
quality Yoko Ogawa 26 0.54 0.46 0.542
quality Ian McEwan 27 0.51 0.49 0.514
quality Ottessa Moshfegh 28 0.51 0.49 0.514
quality Kazuo Ishiguro 29 0.49 0.51 0.486
quality Min Jin Lee 30 0.49 0.51 0.486
style Roxane Gay 1 0.98 0.02 0.986
style Margaret Atwood 2 0.97 0.03 0.972
style Han Kang 3 0.95 0.05 0.958
style Ben Lerner 4 0.94 0.06 0.944
style Chimamanda Ngozi Adichie 5 0.94 0.06 0.944
style Kazuo Ishiguro 6 0.94 0.06 0.944
style Lydia Davis 7 0.94 0.06 0.944
style Marilynne Robinson 8 0.94 0.06 0.944
style Sigrid Nunez 9 0.92 0.08 0.931
style Cheryl Strayed 10 0.91 0.09 0.917
style Junot Diaz 11 0.91 0.09 0.917
style Orhan Pamuk 12 0.91 0.09 0.917
style Min Jin Lee 13 0.86 0.14 0.861
style Sally Rooney 14 0.86 0.14 0.861
style Colson Whitehead 15 0.83 0.17 0.833
style Salman Rushdie 16 0.83 0.17 0.833
style George Saunders 17 0.80 0.20 0.806
style Ian McEwan 18 0.77 0.23 0.778
style Haruki Murakami 19 0.76 0.24 0.764
style Rachel Cusk 20 0.76 0.24 0.764
style Percival Everett 21 0.75 0.25 0.750
style Jhumpa Lahiri 22 0.71 0.29 0.708
style Zadie Smith 23 0.71 0.29 0.708
style Annie Proulx 24 0.66 0.34 0.667
style Jonathan Franzen 25 0.66 0.34 0.667
style Ottessa Moshfegh 26 0.65 0.35 0.653
style Yoko Ogawa 27 0.61 0.39 0.611
style Louise Erdrich 28 0.58 0.42 0.583
style Annie Ernaux 29 0.51 0.49 0.514
style Tony Tulathimutte 30 0.18 0.82 0.181
@@
quality 0.0029 -0.0173 0.0231 0.676 0.580 0.771 0.0031 30
style 0.0089 -0.0194 0.0373 0.760 0.626 0.894 0.0153 30
Non-parametric check. Spearman $\rho=0.052$ ($p=0.791$) for quality and $\rho=-0.040$ ($p=0.839$) for style.
@@
Table LABEL:tab:s8.1 presents the full model examining how AI detectability (Pangram score) influences preferences. The key finding is the significant interaction between pangram_score and setting (quality: $\hat{\beta}\approx 2.90$, $p<0.001$), indicating that fine-tuning attenuates the negative association between detectability and preference observed in in-context prompting. The significant three-way interactions (pangram_score $\times$ setting $\times$ reader_type) further indicate that the detectability penalty and its attenuation differ across reader groups, with college-educated general readers showing less sensitivity to detectability than MFA-trained readers even before fine-tuning.
Table 15: Pangram GLM coefficients with CR2 robust standard errors (clustered by reader). Negative coefficients for pangram_score in in-context prompting indicate detection reduces preference; positive interactions with setting show that fine-tuning attenuates this penalty. Significant three-way interactions indicate the detectability penalty differs across reader groups.
@@
quality 10920 542 (Intercept) 0.990 0.169 5.87 4.43e-09 2.69 1.93 3.74
quality 10920 542 pangram_score -2.010 0.333 -6.03 1.65e-09 0.13 0.07 0.26
quality 10920 542 setting (Fine-tuned) -1.025 0.175 -5.86 4.64e-09 0.36 0.26 0.51
quality 10920 542 reader_type (College-educated) -1.288 0.174 -7.40 1.36e-13 0.28 0.20 0.39
quality 10920 542 pangram_score $\times$ setting 2.904 0.876 3.32 9.13e-04 18.20 3.28 102.00
quality 10920 542 pangram_score $\times$ reader_type 2.613 0.344 7.59 3.24e-14 13.60 6.94 26.80
quality 10920 542 setting $\times$ reader_type 1.313 0.180 7.28 3.34e-13 3.72 2.61 5.29
quality 10920 542 pangram_score $\times$ setting $\times$ reader_type -3.258 0.908 -3.59 3.31e-04 0.04 0.01 0.23
style 10920 535 (Intercept) 0.909 0.146 6.23 4.79e-10 2.48 1.86 3.31
style 10920 535 pangram_score -1.845 0.292 -6.32 2.57e-10 0.16 0.09 0.28
style 10920 535 setting (Fine-tuned) -0.938 0.158 -5.95 2.65e-09 0.39 0.29 0.53
style 10920 535 reader_type (College-educated) -0.925 0.152 -6.10 1.09e-09 0.40 0.29 0.53
style 10920 535 pangram_score $\times$ setting 2.556 0.810 3.15 1.61e-03 12.90 2.63 63.00
style 10920 535 pangram_score $\times$ reader_type 1.878 0.304 6.18 6.26e-10 6.54 3.61 11.90
style 10920 535 setting $\times$ reader_type 0.921 0.163 5.65 1.63e-08 2.51 1.82 3.46
style 10920 535 pangram_score $\times$ setting $\times$ reader_type -1.766 0.844 -2.09 3.64e-02 0.17 0.03 0.89
Notes. CR2 robust standard errors clustered by reader. Readers: 28 MFA-trained and 516 college-educated general readers.
@@
Tables 16 and 17 show model-predicted probabilities at different AI detection levels, corresponding to Figure 2F. For MFA-trained readers evaluating quality, high detection scores (0.9) reduce AI preference to $\sim$31% in in-context prompting but have a point estimate of $\sim$68% after fine-tuning (95% CI 0.370.370.890.89), demonstrating that fine-tuning makes outputs robust to detection-based skepticism. College-educated general readers show a weaker detectability penalty in the in-context condition and a modest positive slope after fine-tuning, with predicted probabilities reaching 0.55–0.67 at high detection scores.
Table 16: Predicted AI preference probabilities for style at different Pangram detection scores. Fine-tuning attenuates the relationship with detection score for MFA-trained readers (with wider uncertainty at high detection). College-educated general readers show a weak relationship with detectability in the in-context condition and a positive slope after fine-tuning.
@@
0.1 In-context Prompting MFA-trained 0.674 0.621 0.722
0.5 In-context Prompting MFA-trained 0.497 0.494 0.500
0.9 In-context Prompting MFA-trained 0.320 0.273 0.372
0.1 Fine-tuned MFA-trained 0.511 0.491 0.530
0.5 Fine-tuned MFA-trained 0.581 0.435 0.714
0.9 Fine-tuned MFA-trained 0.648 0.380 0.847
0.1 In-context Prompting College-educated 0.497 0.481 0.513
0.5 In-context Prompting College-educated 0.500 0.500 0.500
0.9 In-context Prompting College-educated 0.503 0.487 0.520
0.1 Fine-tuned College-educated 0.512 0.505 0.519
0.5 Fine-tuned College-educated 0.593 0.544 0.641
0.9 Fine-tuned College-educated 0.670 0.581 0.748
Table 17: Predicted AI preference probabilities for quality. MFA-trained readers show strong detection sensitivity in in-context prompting ($\sim$0.69 $\to$ $\sim$0.31) that attenuates after fine-tuning (point estimate at 0.9 $\sim$0.68; 95% CI 0.37–0.89). College-educated general readers show a weaker in-context penalty and modest positive slopes after fine-tuning.
@@
0.1 In-context Prompting MFA-trained 0.688 0.628 0.742
0.5 In-context Prompting MFA-trained 0.496 0.493 0.499
0.9 In-context Prompting MFA-trained 0.306 0.254 0.363
0.1 Fine-tuned MFA-trained 0.514 0.490 0.537
0.5 Fine-tuned MFA-trained 0.602 0.431 0.751
0.9 Fine-tuned MFA-trained 0.683 0.373 0.887
0.1 In-context Prompting College-educated 0.441 0.424 0.457
0.5 In-context Prompting College-educated 0.501 0.500 0.501
0.9 In-context Prompting College-educated 0.561 0.544 0.578
0.1 Fine-tuned College-educated 0.504 0.497 0.510
0.5 Fine-tuned College-educated 0.529 0.479 0.578
0.9 Fine-tuned College-educated 0.553 0.461 0.643
@@
Table 18 shows that cliché density has the strongest correlation with AI detection (Pearson $r=0.60$), while readability ease is negatively correlated (r=0.23r=-0.23), suggesting AI text is marked by formulaic phrases but simpler syntax. Among MFA-trained readers, where the detectability penalty is concentrated, approximately 16% of the detection effect on preference is mediated through cliché density in the in-context prompting condition; after fine-tuning, this mediation drops to a statistically insignificant 1.3%, indicating that fine-tuning substantially reduces rather than merely masks these stylistic signatures. Because college-educated general readers did not penalize AI-detectable text in either condition, mediation analysis does not apply to this group.
Table 18: Correlations between Pangram AI detection scores and stylometric features. These correlations are computed at the excerpt level ($N=330$ unique excerpts) and are unchanged by the expanded reader sample.
@@
Table 19: Cost comparison for generating 100,000 words. Fine-tuning plus inference costs (\$25–\$276) represent about 0.1% to 1.1% of professional writer compensation (\$25,000), with a median of approximately 0.3%.
@@
Note. Inference cost of \$3 per 100k words follows Figure 4c; per-token assumptions are documented in the repository.
@@
    • MFA-trained readers: 28 (planned:  25)
    • College-educated general readers: 516 (planned: 120)